If f = (1, 2), (2, 3), (3, 4), (4, 5),
g = (1, -2), (3, -3), (5, -5), and
h = (1, 0), (2, 1), (3, 2),

find the following and state the domain:

f / h
1
Expert's answer
2013-01-15T09:11:08-0500
Determine value of x where both function (f, h) have value:
x = 1 -> f(1) = 2 and h(1) = 0
x = 2 -> f(2) = 3 and h(2) = 1
x = 3 -> f(3) = 4 and h(3) = 2

Find function f/h when x = 1, 2, 3:
f/h(1) = f(1)/h(1) = 2/0 & undefined
f/h(2) = f(2)/h(2) = 3/1 = 3
f/h(3) = f(3)/h(3) = 4/2 = 2

So f/h& = (2, 3), (3, 2)

domain of a function is the set of "input" or argument values (x) for which the function is defined, so

domain: {2, 3} &